OpenStudy (anonymous):

what are the steps to finding the slope-intercept if you have 2x-3y=12?

OpenStudy (anonymous):

Add 2, divide by 3

OpenStudy (anonymous):

Basically you have to solve for y. 2x - 3y = 12 (subtract 2x from both sides) -3y = -2x + 12 (now divide by -3 to get y by itself) y = (-2/-3)x + 12/-3 y = 2/3x - 4 -- slope intercept form any questions ?